# RTMO: Towards High-Performance One-Stage Real-Time Multi-Person Pose Estimation

### 12 Dec 2023 v1

@@ Abstract
Real-time multi-person pose estimation presents significant challenges in balancing speed and precision. While two-stage top-down methods slow down as the number of people in the image increases, existing one-stage methods often fail to simultaneously deliver high accuracy and real-time performance. This paper introduces RTMO, a one-stage pose estimation framework that seamlessly integrates coordinate classification by representing keypoints using dual 1-D heatmaps within the YOLO architecture, achieving accuracy comparable to top-down methods while maintaining high speed. We propose a dynamic coordinate classifier and a tailored loss function for heatmap learning, specifically designed to address the incompatibilities between coordinate classification and dense prediction models. RTMO outperforms state-of-the-art one-stage pose estimators, achieving 1.1% higher AP on COCO while operating about 9 times faster with the same backbone. Our largest model, RTMO-l, attains 74.8% AP on COCO val2017 and 141 FPS on a single V100 GPU, demonstrating its efficiency and accuracy. The code and models are available at https://github.com/open-mmlab/mmpose/tree/dev-1.x/projects/rtmo.
@@ 1 Introduction
Consequently, RTMO achieves accuracy comparable to real-time top-down methods and exceeds that of other lightweight one-stage methods, as shown in Figure 1. Moreover, RTMO demonstrates superior speed when processing multiple instances in an image, outpacing top-down methods with similar accuracy. Notably, the RTMO-l model attains a 74.8% Average Precision (AP) on the COCO val2017 dataset [23] and exhibited 141 frames per second (FPS) on the NVIDIA V100 GPU. On the CrowdPose benchmark [18], RTMO-l achieves 73.2% AP, a new state of the art for one-stage methods. The key contributions of this work include:
An innovative coordinate classification technique tailored for dense prediction scenarios, utilizing coordinate bins for precise keypoint localization while addressing the challenges posed by the varying sizes and complexities of instances.
A new real-time one-stage MPPE approach that seamlessly integrates coordinate classification with the YOLO architecture, achieving an optimal balance of performance and speed among existing top-down and one-stage MPPE methods.
@@ 2.1 One-Stage Pose Estimator
Inspired by advancements in one-stage object detection algorithms [40, 24, 9, 51, 7], a series of one-stage pose estimation methods have emerged [39, 34, 51, 10, 30]. These methods execute MPPE within a single forward pass and directly regress instance-specific keypoints from predetermined root locations. Alternative approaches such as PETR [37] and ED-Pose [46] treat pose estimation as a set prediction problem, establishing a fully end-to-end framework for keypoint regression. Beyond regression-based solutions, techniques like FCPose [31], InsPose [35] and CID [42] utilize dynamic convolution or attention mechanisms to generate instance-specific heatmaps for keypoint localization.
Compared to two-stage pose estimation methods, one-stage approaches eliminate the need for pre-processing (e.g., human detection for top-down methods) and post-processing (e.g., keypoint grouping for bottom-up methods). This results in two benefits: 1) consistent inference time, irrespective of the number of instances in the image; and 2) a simplified pipeline that facilitates deployment and practical use. Despite these advantages, the existing one-stage methods struggle to balance high accuracy with real-time inference. High accuracy models [42, 46] often depend on resource-intensive backbones like HRNet [38] or Swin [25], making real-time estimation challenging. Conversely, real-time models [33, 30] compromise on performance. Our model addresses this trade-off, delivering both high accuracy and fast real-time inference.
@@ 2.2 Coordinate Classification
SimCC [22] and RTMPose [15] have adopted coordinate classification for pose estimation. These methods classify keypoints based on sub-pixel bins along horizontal and vertical axes, achieving spatial discrimination without relying on high-resolution feature maps. This balances accuracy and speed effectively. However, spanning bins across the entire image for dense prediction methods is impractical due to the need for a vast number of bins to diminish quantization error, which leads to inefficiency from many bins being superfluous for individual instances. DFL [20] sets bins around a predefined range near each anchor, which may not encompass the keypoints of large instances and could induce significant quantization errors for small instances. Our approach assigns bins based on each instance’s size, placing them within localized regions, which optimizes bin utilization, ensures coverage of keypoints and minimizes quantization error.
@@ 3 Methodology
In our model, we adopt a YOLO-like architecture, illustrated in Figure 2. The backbone is CSPDarknet [9], and we process the last three feature maps using a Hybrid Encoder [28], yielding spatial features and with respective downsampling rates of 16 and 32. Each pixel in these features maps to a grid cell uniformly distributed on the original image plane. The network head, utilizing dual convolution blocks at each spatial level, generates a score and corresponding pose features for every grid cell. These pose features are utilized to predict bounding boxes, keypoint coordinates, and visibility. The generation of 1-D heatmap predictions through the Dynamic Coordinate Classifier is detailed in Sec. 3.1 while the proposed heatmap loss based on MLE is presented in Sec. 3.2. The complete training and inference procedures are outlined in Sec. Sec. 3.3.
@@ 3.1 Dynamic Coordinate Classifier
Coordinate classification technique employed in top-down pose estimators allocates bins across the entire input image [22, 15]. This approach leads to bin wastage in one-stage methods, as each subject occupies only a small portion of the image. DFL [20] sets bins within a predefined range near each anchor, which may miss keypoints in larger instances and cause significant quantization errors in smaller ones. DCC addresses these limitations by dynamically assigning bins to align with each instance’s bounding box, ensuring localized coverage. The bounding boxes are initially regressed using a pointwise convolution layer and then expanded by a factor of 1.25 to cover all keypoints, even in cases of inaccurate predictions. These expanded bounding boxes are uniformly divided into bins along horizontal and vertical axes. The x-coordinate for each horizontal bin is calculated using:
where are the left and right sides of the bounding box, and index varies from 1 to . The y-axis bins are computed similarly.
@@ 3.2 MLE for Coordinate Classification
Unlike KLD, our MLE loss allows for learnable variance, representing uncertainty. This uncertainty learning framework automatically adjusts the difficulty of various samples [12, 5]. For hard samples, the model predicts a large variance to ease optimization. For simple samples, it predicts a small variance, aiding in accuracy improvement. With KLD, adopting a learnable variance is problematic - the model leans to predict a large variance to flatten the target distribution as this simplifies learning.
@@ 3.3 Training and Inference
Our model, adhering to a YOLO-like structure, employs dense grid prediction for human detection and pose estimation. It is crucial for the model to differentiate between positive and negative grids. We extend SimOTA [9] for training, assigning positive grids based on grid scores, bounding box regression and pose estimation accuracy. The head’s score branch classifies these grids, supervised by varifocal loss [50] , with target scores being the Object Keypoint Similarity (OKS) between the predicted pose and the assigned ground truth for each grid.
@@ 4.1 Settings
Experiments were primarily conducted on the COCO2017 Keypoint Detection benchmark [23], comprising approximately 250K person instances with 17 keypoints. Performance comparisons were made with state-of-the-art methods on both the val2017 and test-dev sets. To explore our model’s performance ceiling, training was also extended to include additional datasets: CrowdPose [18], AIC [43], MPII [1], JHMDB [14], Halpe [8], and PoseTrack18 [2]. These annotations were converted to COCO format. RTMO was further evaluated on the CrowdPose benchmark [18] which is known for its high complexity due to crowded and occluded scenes, comprising 20K images and approximately 80K persons with 14 keypoints. OKS-based Average Precision (AP) served as the evaluation metric for both datasets.
@@
During training, we adopt the image augmentation pipeline from YOLOX [9], incorporating mosaic augmentation, random color adjustments, geometric transformations, and MixUp [49]. Training images are resized to dimensions [480, 800]. Epoch counts are set to 600 and 700 for the COCO and CrowdPose datasets, respectively. The training process is divided into two stages: the first involves training both the proxy branch and DCC using pose annotations, and the second shifts the target of the proxy branch to the decoded pose from DCC. The AdamW optimizer [26] is used with a weight decay of 0.05, and training is performed on Nvidia GeForce RTX 3090 GPUs with batch size 256. Initial learning rates are set to and for the two training phases, decaying to via cosine annealing. For inference, images are resized to 640. CPU latency is measured on an Intel Xeon Gold CPU using ONNXRuntime. GPU latency is tested on an NVIDIA V100 GPU with ONNXRuntime and with TensorRT using half-precision floating-point (FP16) format.
@@ 4.2 Benchmark Results
Our evaluation of RTMO against leading one-stage pose estimators on the COCO test-dev is presented in Table 1. RTMO showcases significant advancements in speed and precision. Specifically, RTMO-s outperforms PETR [36] using a ResNet-50 [11] backbone, being ten times quicker yet maintaining similar accuracy. Compared to lightweight models like KAPAO and YOLO-Pose, RTMO consistently outperforms in accuracy across different model sizes. When trained on COCO train2017, RTMO-l has the second-best performance among all tested models. The highest-performing model, ED-Pose [46] with a Swin-L [25] backbone, is quite heavy and not deployment-friendly. RTMO, using the same ResNet-50 backbone, surpassed ED-Pose by 1.1% in AP and was faster. Additionally, transferring ED-Pose to the ONNX format resulted in a slower latency than its PyTorch model, about 1.5 seconds per frame. By contrast, the ONNX model of RTMO-l processes an image in just 19.1ms. With further training on additional human pose datasets, RTMO-l performs best among one-stage pose estimators in terms of accuracy.
@@
4.3 Quantitative Results
RTMO employs the coordinate classification method and shows strong performance in challenging multi-person scenarios, where individuals are relatively small and frequent occlusions occur. Figure 4 reveals that RTMO generates spatially accurate heatmaps even under these difficult conditions, facilitating robust and context-aware predictions for each keypoint.
@@ 4.4 Ablation Study
To assess the effectiveness of coordinate classification against regression, we replaced the model’s 1-D heatmap generation with a fully-connected layer for regression, supervised by the OKS loss [30]. Table 3 compares the performances. Using the DCC module and MLE loss, coordinate classification outperforms regression with 2.1% AP on the COCO.
@@
Contrary to other pose estimation methods with coordinate classification that use KLD as the loss function, our research indicates its inadequacy for RTMO. Evidence from Table 3 demonstrates lower accuracy with KLD compared to our MLE loss. We attribute this improvement to the inclusion of learnable variance in the MLE loss function, which helps to balance the learning between hard and easy samples. Specifically, in a one-stage pose estimator framework, the difficulty level varies for each grid and is influenced by several factors, including the assigned instance’s pose and size, as well as the relative position between the grid and instance. KLD fails to account for this variability, making it less effective in this context.
@@
We firstly adopt a static coordinate classification strategy similar to DFL [20] where bins are distributed in a fixed range around each grid. This approach outperformed the regression method on the COCO dataset but underperformed on CrowdPose. Introducing the Dynamic Bin Allocation (DBA) strategy to this baseline resulted in decreased performance on both datasets. This is reasonable, as the semantic of each bin vary across different samples without corresponding representation adjustments. This issue was rectified by incorporating Dynamic Bin Encoding (DBE). With DBE, our DCC methods exceeded the efficacy of the static strategy on both datasets. Furthermore, Without dynamic bin encoding (DBE), the probabilities of nearby bins can vary significantly, as shown in Figure 5, contradicting the expectation that adjacent spatial locations should have similar probabilities. In contrast, incorporating DBE leads to smoother output heatmaps, indicating improved decoder training by enabling representation vectors that better capture similarities between nearby locations.
@@
| {P4, P5} | 125.4 | 19.09 | 72.4 | 76.2 | |
Feature pyramids [17] leverage multi-scale features for detecting instances of varying sizes; deeper features typically detect larger objects. Our initial model, inspired by YOLOX, used P3, P4, P5 features with strides of 8, 16, and 32 pixels. However, P3 contributed 78.5% of the FLOPs in the model head while accounting for 10.7% of correct detections. To improve efficiency, we focused on P4, P5. As shown in Table 4, omitting P3 led to significant speed gains with minimal accuracy loss, indicating that P4 and P5 alone are effective for multi-person pose estimation. This demonstrates that the role of P3 in detecting smaller instances can be sufficiently compensated by the remaining features.

### 08 Apr 2024 v2

@@ Abstract
Real-time multi-person pose estimation presents significant challenges in balancing speed and precision. While two-stage top-down methods slow down as the number of people in the image increases, existing one-stage methods often fail to simultaneously deliver high accuracy and real-time performance. This paper introduces RTMO, a one-stage pose estimation framework that seamlessly integrates coordinate classification by representing keypoints using dual 1-D heatmaps within the YOLO architecture, achieving accuracy comparable to top-down methods while maintaining high speed. We propose a dynamic coordinate classifier and a tailored loss function for heatmap learning, specifically designed to address the incompatibilities between coordinate classification and dense prediction models. RTMO outperforms state-of-the-art one-stage pose estimators, achieving 1.1% higher AP on COCO while operating about 9 times faster with the same backbone. Our largest model, RTMO-l, attains 74.8% AP on COCO val2017 and 141 FPS on a single V100 GPU, demonstrating its efficiency and accuracy. The code and models are available at https://github.com/open-mmlab/mmpose/tree/main/projects/rtmo.
@@ 1 Introduction
Consequently, RTMO achieves accuracy comparable to real-time top-down methods and exceeds other lightweight one-stage methods, as shown in Figure 1. Moreover, RTMO demonstrates superior speed when processing multiple instances in an image, outpacing top-down methods with similar accuracy. Notably, RTMO-l attains a 74.8% Average Precision (AP) on the COCO val2017 dataset [24] and exhibited 141 frames per second (FPS) on the NVIDIA V100 GPU. On the CrowdPose benchmark [19], RTMO-l achieves 73.2% AP, a new state of the art for one-stage methods. The key contributions of this work include:
An innovative coordinate classification technique tailored for dense prediction, utilizing coordinate bins for precise keypoint localization while addressing varying instance sizes and complexities.
A new real-time one-stage MPPE approach that seamlessly integrates coordinate classification with the YOLO architecture, achieving an optimal balance of performance and speed among existing MPPE methods.
@@ 2.1 One-Stage Pose Estimator
Inspired by advancements in one-stage object detection algorithms [41, 25, 10, 52, 8], a series of one-stage pose estimation methods have emerged [40, 35, 52, 11, 31]. These methods perform MPPE in a single forward pass and directly regress instance-specific keypoints from predetermined root locations. Alternative approaches such as PETR [38] and ED-Pose [47] treat pose estimation as a set prediction problem for end-to-end keypoint regression. Beyond regression-based solutions, techniques like FCPose [32], InsPose [36], and CID [43] utilize dynamic convolution or attention mechanisms to generate instance-specific heatmaps for keypoint localization.
Compared to two-stage pose estimation methods, one-stage approaches eliminate the need for pre-processing (e.g., human detection for top-down methods) and post-processing (e.g., keypoint grouping for bottom-up methods). This results in two benefits: 1) consistent inference time, irrespective of the number of instances in the image; and 2) a simplified pipeline that facilitates deployment and practical use. Despite these advantages, the existing one-stage methods struggle to balance high accuracy with real-time inference. High-accuracy models [43, 47] often depend on resource-intensive backbones like HRNet [39] or Swin [26], making real-time estimation challenging. Conversely, real-time models [34, 31] compromise on performance. Our model addresses this trade-off, delivering both high accuracy and fast real-time inference.
@@ 2.2 Coordinate Classification
SimCC [23] and RTMPose [16] have adopted coordinate classification for pose estimation, classifying keypoints into sub-pixel bins along horizontal and vertical axes to achieve spatial discrimination without high-resolution features, balancing accuracy and speed. However, spanning bins across the entire image for dense prediction methods is impractical due to the vast number of bins needed to diminish quantization error, which leads to inefficiency from many bins being superfluous for individual instances. DFL [21] sets bins around a predefined range near each anchor, which may not cover the keypoints of large instances and could induce significant quantization errors for small instances. Our approach assigns bins within localized regions scaled to each instance’s size, which optimizes bin utilization, ensures keypoint coverage, and minimizes quantization error.
@@ 3 Methodology
In our model, we adopt a YOLO-like architecture, illustrated in Figure 2. The backbone is CSPDarknet [10], and we process the last three feature maps using a Hybrid Encoder [29], yielding spatial features and with respective downsampling rates of 16 and 32. Each pixel in these features maps to a grid cell uniformly distributed on the original image plane. The network head, utilizing dual convolution blocks at each spatial level, generates a score and corresponding pose features for every grid cell. These pose features are utilized to predict bounding boxes, keypoint coordinates, and visibility. The generation of 1-D heatmap predictions through the Dynamic Coordinate Classifier is detailed in Sec. 3.1 while the proposed heatmap loss based on MLE is presented in Sec. 3.2. The complete training and inference procedures are outlined in Sec. 3.3.
@@ 3.1 Dynamic Coordinate Classifier
Coordinate classification technique employed in top-down pose estimators allocates bins across the entire input image [23, 16], leading to bin wastage in one-stage methods since each subject occupies only a small portion. DFL [21] sets bins within a predefined range near each anchor, which may miss keypoints in larger instances and cause significant quantization errors in smaller ones. DCC addresses these limitations by dynamically assigning bins to align with each instance’s bounding box, ensuring localized coverage. The bounding boxes are initially regressed using a pointwise convolution layer and then expanded by 1.25x to cover all keypoints, even in cases of inaccurate predictions. These expanded bounding boxes are uniformly divided into bins along horizontal and vertical axes. The x-coordinate for each horizontal bin is calculated using:
where are the left and right sides of the bounding box, and index varies from 1 to . The y-axis bins are computed similarly. We empirically use and for all models.
@@ 3.2 MLE for Coordinate Classification
Unlike KLD, our MLE loss allows for learnable variance, representing uncertainty. This uncertainty learning framework automatically adjusts the difficulty of various samples [13, 5]. For hard samples, the model predicts a large variance to ease optimization. For simple samples, it predicts a small variance, aiding in accuracy improvement. With KLD, adopting a learnable variance is problematic - the model leans to predict a large variance to flatten the target distribution as this simplifies learning. More discussion can be found in Sec. 4.4.
@@ 3.3 Training and Inference
Our model, adhering to a YOLO-like structure, employs dense grid prediction for human detection and pose estimation. It is crucial for the model to differentiate between positive and negative grids. We extend SimOTA [10] for training, assigning positive grids based on grid scores, bounding box regression, and pose estimation accuracy. The head’s score branch classifies these grids, supervised by varifocal loss [51] , with target scores being the Object Keypoint Similarity (OKS) between the predicted pose and the assigned ground truth for each grid.
@@ 4.1 Settings
Experiments were primarily conducted on the COCO2017 Keypoint Detection benchmark [24], comprising approximately 250K person instances with 17 keypoints. Performance comparisons were made with state-of-the-art methods on both the val2017 and test-dev sets. To explore our model’s performance ceiling, training was also extended to include additional datasets: CrowdPose [19], AIC [44], MPII [1], JHMDB [15], Halpe [9], and PoseTrack18 [2]. These annotations were converted to COCO format. RTMO was further evaluated on the CrowdPose benchmark [19], which is known for its high complexity due to crowded and occluded scenes, comprising 20K images and approximately 80K persons with 14 key points. OKS-based Average Precision (AP) served as the evaluation metric for both datasets.
@@
During training, we adopt the image augmentation pipeline from YOLOX [10], incorporating mosaic augmentation, random color adjustments, geometric transformations, and MixUp [50]. Training images are resized to dimensions [480, 800]. Epoch counts are set to 600 and 700 for the COCO and CrowdPose datasets, respectively. The training process is divided into two stages: the first involves training both the proxy branch and DCC using pose annotations, and the second shifts the target of the proxy branch to the decoded pose from DCC. The AdamW optimizer [27] is used with a weight decay of 0.05, and training is performed on Nvidia GeForce RTX 3090 GPUs with batch size 256. Initial learning rates are set to and for the two training phases, decaying to via cosine annealing. For inference, images are resized to 640. CPU latency is measured on an Intel Xeon Gold CPU using ONNXRuntime. GPU latency is tested on an NVIDIA V100 GPU with ONNXRuntime and with TensorRT using half-precision floating-point (FP16) format. MMPose [7] toolbox is used to implement RTMO models.
@@ 4.2 Benchmark Results
Our evaluation of RTMO against leading one-stage pose estimators on the COCO test-dev is presented in Table 1. RTMO showcases significant advancements in speed and precision. Specifically, RTMO-s outperforms PETR [37] using a ResNet-50 [12] backbone, being 10x faster while maintaining similar accuracy. Compared to lightweight models like KAPAO and YOLO-Pose, RTMO consistently outperforms in accuracy across different model sizes. When trained on COCO train2017, RTMO-l has the second-best performance among all tested models. The highest-performing model, ED-Pose [47] with a Swin-L [26] backbone, is quite heavy and not deployment-friendly. RTMO, using the same ResNet-50 backbone, surpassed ED-Pose by 1.1% in AP and was faster. Additionally, transferring ED-Pose to the ONNX format resulted in a higher latency than its PyTorch model, about 1.5 seconds per frame. By contrast, the ONNX model of RTMO-l processes an image in just 19.1ms. With further training on additional human pose datasets, RTMO-l performs best among one-stage pose estimators in terms of accuracy.
@@
4.3 Qualitative Results
RTMO utilizes coordinate classification and demonstrates strong performance in challenging multi-person scenarios with small individuals and frequent occlusions. Figure 4 reveals that RTMO generates spatially accurate heatmaps even under these difficult conditions, facilitating robust and context-aware predictions for each keypoint.
@@ 4.4 Ablation Study
To assess the effectiveness of coordinate classification against regression, we replaced the model’s 1-D heatmap generation with a fully connected layer for regression, supervised by the OKS loss [31]. Table 3 compares the performances. Using the DCC module and MLE loss, coordinate classification outperforms regression with 2.0% AP on the COCO.
@@
Compare to other pose estimation methods with coordinate classification that use KLD loss, our research indicates its inadequacy for RTMO. Table 3 demonstrates our MLE loss achieves higher accuracy than KLD. This improvement stems from the learnable variance in the MLE loss, which helps to balance the learning between hard and easy samples. In a one-stage pose estimator, the difficulty varies per grid due to factors like instance pose, size, and relative grid position, as visualized in Figure 5. Grids with higher OKS (easier) have lower variance in MLE loss, and vice versa. KLD fails to account for this variability, making it less effective in this context.
@@
We firstly adopt a static coordinate classification strategy similar to DFL [21] where bins are distributed in a fixed range around each grid. This approach outperformed the regression method on the COCO dataset but underperformed on CrowdPose. Introducing the Dynamic Bin Allocation (DBA) strategy to this baseline resulted in decreased performance on both datasets. This is reasonable, as the semantics of each bin vary across different samples without corresponding representation adjustments. This issue was rectified by incorporating Dynamic Bin Encoding (DBE). With DBE, our DCC methods exceeded the efficacy of the static strategy on both datasets. Furthermore, Without dynamic bin encoding (DBE), the probabilities of nearby bins can vary significantly, as shown in Figure 6, contradicting the expectation that adjacent spatial locations should have similar probabilities. In contrast, incorporating DBE leads to smoother output heatmaps, indicating improved decoder training by enabling representation vectors that better capture similarities between nearby locations.
@@
| {P4, P5} | 125.4 | 19.09 | 72.5 | 76.6 | |
Feature pyramids [18] leverage multi-scale features for detecting instances of varying sizes; deeper features typically detect larger objects. Our initial model used P3, P4, P5 features with strides of 8, 16, and 32 pixels. However, P3 contributed 78.5% of the FLOPs in the model head while accounting for 10.7% of correct detections. To improve efficiency, we focused on P4, P5. As shown in Table 4, omitting P3 led to significant speed gains with minimal accuracy loss, indicating that P4 and P5 alone are effective for multi-person pose estimation. This suggests that the role of P3 in detecting smaller instances can be compensated by the remaining features.
@@ 5 Conclusion
Acknowledgments
We thank the reviewers for their helpful comments. This work was supported by the National Key R&D Program of China (No. 2022ZD0161600) and the Special Foundations for the Development of Strategic Emerging Industries of Shenzhen (Nos. JCYJ20200109143035495 & CJGJZD20210408092804011)